OpenStudy (anonymous):

Find two positive numbers whose product is 225 and whose sum is a minimum

OpenStudy (anonymous):

-15^=225 -15+-15=-30

OpenStudy (anonymous):

sooo what are the 2 numbers?

OpenStudy (anonymous):

you cant have two numberwhose sum is negative and product is positive other than a square number ie, 4^2,5^2 only way is if you have -25(-1)(9)=225 -25+-1+9<0

OpenStudy (anonymous):

Let S = x + y. Since xy = 225, y = 225/x. Substitute into the sum to get S = x + 225/x. Then S' = 1 - 225/x^2, implying that 15/x = 1. Then x = 15. The other number is also 15.

OpenStudy (anonymous):

hell yea it is thanks man

OpenStudy (anonymous):

:^)